OpenStudy (anonymous):

Write an expression for the slope of a line whose equation is Ax+By+C=0

OpenStudy (anonymous):

By = - Ax - C A C y = - --- x - ---- B B so slope = -A/B

OpenStudy (anonymous):

Hey, you do know there is an equation editor, right? It would have made typing that much easier.

OpenStudy (anonymous):

For fractions it's not quite as clear when you use the editor so I can see why some folks type it as above.

OpenStudy (anonymous):

You just need to look up the LaTex for them. It's pretty much LaTex compatible.

OpenStudy (anonymous):

Well let me test it since I'm curious: \[ \frac{1}{2}\]

OpenStudy (anonymous):

Guess you're right :) They need to embed that in the GUI for the site.

OpenStudy (anonymous):

Looks like the form is \frac{nominator}{denominator}. Thanks.

OpenStudy (anonymous):

*It's Numerator

OpenStudy (anonymous):

That it is :)